# Differentially private and decentralized randomized power method

### v2

@@ Abstract
The randomized power method has gained significant interest due to its simplicity and efficient handling of large-scale spectral analysis and recommendation tasks. However, its application to large datasets containing personal information (e.g. web interactions, search history, personal tastes) raises critical privacy concerns. This paper addresses these issues by proposing enhanced privacy-preserving variants of the method, which reduce the variance of the noise introduced to achieve Differential Privacy (DP). We also adapt the method to a decentralized framework with a low computational and communication overhead, while preserving the accuracy. We leverage Secure Aggregation to run the algorithm on data distributed among multiple users or devices, while using a noise scale in the decentralized setting that is similar to the one in the centralized setting yet without revealing individual data. We improve upon existing convergence bounds for both the centralized and decentralized versions.
Machine Learning, ICML
@@ 1 Introduction
The randomized power method is a simple and computationally efficient tool for solving large-scale linear algebra problems, particularly in spectral analysis and recommendation systems. It approximates the principal singular vectors of a matrix with low computational complexity (nearly linear), making it particularly suitable for analyzing large modern datasets.
@@
In large-scale machine learning systems, protecting user privacy is paramount. Datasets often contain sensitive personal or organizational information, so privacy has become a critical requirement. The standard randomized power method does not inherently provide privacy guarantees. While its output might seem less sensitive than the input data, there is no formal guarantee that private information embedded in the data cannot be inferred.
@@
To address and quantify privacy leakages, Differential Privacy (DP) has emerged as a powerful mathematical framework that provides strong formal guarantees and mitigate potential privacy leaks of an algorithm. DP ensures that the output of an algorithm reveals little about any individual record in the input. Several works have attempted to apply DP to the randomized power method. For instance, Hardt & Price (2014) and Balcan et al. (2016) introduced differentially private versions of the method, and Guo et al. (2021); Wang & Xu (2020) explored federated approaches to private power iteration. Other close works explore centralized differentially private versions of PCA (Liu et al., 2022) and federated (Wang & Xu, 2020; Briguglio et al., 2023) and give optimal convergence bounds given certain assumptions.
@@
Despite these advancements, existing approaches suffer from several limitations. First, their performance heavily depends on the number of singular vectors being computed (Hardt & Price, 2014; Balcan et al., 2016; Guo et al., 2021; Liu et al., 2022; Wang & Xu, 2020), which impacts both utility and privacy guarantees. Second, they are primarily designed for centralized settings (Hardt & Price, 2014; Liu et al., 2022), where a trusted curator is assumed to hold the data. Moreover, some methods (Liu et al., 2022) make strong assumptions about the underlying data distribution (i.e. that the data is sub-Gaussian) which makes it harder to use these methods in practice. Some federated versions (Briguglio et al., 2023) claim to guarantee privacy by use of the federated setting, although it has been shown that decentralization does not offer privacy by design (Geiping et al., 2020). Finally, no fully decentralized versions exist to our knowledge. The previously proposed federated versions (Balcan et al., 2016; Guo et al., 2021; Wang & Xu, 2020) use public channel communication and ignore advances in Multi-Party-Computation (MPC). This makes them unsuitable for decentralized environments, such as distributed recommender systems and social networks, where data is partitioned across users and/or devices and communications are restricted to a predefined communication graph.
@@ 1.1 Contributions
The core focus of this article is to introduce improved privacy-preserving randomized power methods and to extend them to a decentralized setting, while maintaining computational efficiency and performance. We propose new convergence bounds for the (decentralized) privacy-preserving power method; these show that our proposed algorithm has a weaker dependence on the number of singular vectors that are computed.
@@
We propose an improved differentially private version of the randomized power method, with tighter convergence bounds, reducing the noise required to achieve privacy and therefore improving the accuracy of the method. In the process of deriving these bounds, we present a new, more straightforward proof of a privacy result for the differentially private randomized power method. [^1] This proof addresses some minor mistakes in a privacy proof from Hardt & Roth (2012, 2013) that have been reproduced in several follow-up works (Hardt & Price, 2014; Balcan et al., 2016). The modified result allows us to use a wider range of privacy parameters.
Finally, we propose decentralized versions of the algorithms. These are particularly relevant in contexts where data is distributed across multiple users or devices, as is often the case in recommender systems. To address the privacy concerns of the users, without forcing them to rely on a trusted curator, we propose a federated privacy-preserving version of the method using Secure Aggregation (Bell et al., 2020), a form of Multi-Party Computation (MPC) and a fully decentralized version using GOPA (GOssip noise for Private Averaging) (Sabater et al., 2022). This allows users to keep their data local while contributing to a privacy-preserving global computation. Our decentralized approach preserves the effectiveness of the centralized version of the randomized power method while incorporating the privacy advantages of local differential privacy (DP). Notably, by employing Secure Aggregation, we are able to set the noise scale comparable to that of central DP, enhancing accuracy. This makes the proposed method a promising candidate for decentralized and privacy-conscious applications.
2 Background Material
2.1 Notation:
Let $\bm{A}\in\mathbb{R}^{n\times n}$ be a real-valued positive semi-definite matrix, and let $k$ denote the target rank, where typically $k\ll n$. We also define $\eta$ as a small positive scalar that represents the precision of the approximation.
The eigenvalue decomposition of $\bm{A}$ is given by:
$\bm{A}=\bm{U}\bm{\Lambda}\bm{U}^{\top},$ (1)
where $\bm{U}$ is the matrix of eigenvectors of $\bm{A}$, and $\bm{\Lambda}$ is a diagonal matrix containing the corresponding eigenvalues on its diagonal.
Let $\mathcal{N}(\mu,\sigma^{2})^{n\times p}$ be a ($n\times p$) random matrix of which each element is a random variable independent and identically distributed (i.i.d.) according to a Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$.
@@ 2.1 Notation:
For any matrix $\bm{X}$, let its element-wise maximum norm be $\|\bm{X}\|_{\max}=\max_{i,j}\|\bm{X}_{ij}\|$, where $\bm{X}_{ij}$ is the $(i,j)$-th entry of matrix $\bm{X}$. We denote by $\bm{X}_{j:}$ the $j$-th row of $\bm{X}$.
We also define two coherence measures for the matrix $\bm{A}$:
@@
  • The $\mu_{1}$-coherence of $\bm{A}$ is the maximum row sum of the absolute values of its eigenvector entries, defined as $\mu_{1}(\bm{A})=\|\bm{U}\|_{\infty}=\max_{i}(\sum_{j}|\bm{U}_{ij}|)$.
Finally, we will also use the QR decomposition of matrix obtained using the Gram-Schmidt procedure. Given a matrix $\bm{X}\in\mathbb{R}^{n\times p}$, the QR decomposition factorizes it as $\bm{X}=\bm{Q}\bm{R}$, where $\bm{Q}$ is an orthonormal matrix (i.e., $\bm{Q}^{\top}\bm{Q}=\bm{I}$) and $\bm{R}$ is an upper triangular matrix.
2.2 Differential privacy:
2.2.1 Approximate differential privacy:
A randomized algorithm $\mathcal{M}$ provides ($\epsilon,\delta$)-differential privacy if for all datasets $D_{1}$ and $D_{2}$ differing on at most one element (adjacent datasets), and for all measurable subsets $S\subseteq\text{Range}(\mathcal{M})$, the following holds:
@@ 2.2.1 Approximate differential privacy:
where $\epsilon$ is a small positive parameter representing the privacy loss (smaller values indicate stronger privacy guarantees), and $\delta$ is a small probability representing the chance that the privacy guarantee may fail, allowing for a small probability of privacy breach.
2.2.2 $\ell_{2}$-sensitivity:
The $\ell_{2}$-sensitivity of a function $f$, denoted as $\Delta_{2}(f)$ or $\Delta_{2}$ for brevity, measures the maximum change in the output of $f$ when its input changes by at most one element (for neighboring datasets $D_{1}\sim D_{2}$). It is defined as:
$\Delta_{2}=\max_{D_{1}\sim D_{2}}\|f(D_{1})-f(D_{2})\|_{2},$ (3)
where $\|\cdot\|_{2}$ represents the $\ell_{2}$-norm.
2.2.3 Zero-Concentrated Differential Privacy (zCDP):
$D_{\alpha}(\mathcal{M}(D_{1})\|\mathcal{M}(D_{2}))\leq\rho\alpha\,,$ (4)
The following lemma, introduced by Bun & Steinke (2016), specifies how the addition of Gaussian noise can be used to construct a randomized algorithm that satisfies zCDP.
Lemma 2.1 (Gaussian Mechanism (Proposition 1.6 (Bun & Steinke, 2016))).
The next lemma, which is a generalized version of a result presented by Bun & Steinke (2016), explains how a randomized algorithm, constructed by recursively composing a sequence of zCDP-satisying randomized algorithms, also satisfies zCDP.
Lemma 2.2 (Adaptive composition (Generalization from Lemma 2.3 (Bun & Steinke, 2016))).
Then $M^{\prime\prime}$ satisfies $(\sum_{i=1}^{L}\rho_{i})$-zCDP.
2.3 Privacy preserving power iteration:
@@
Let $k$ denote the target rank, $q$ a small real and $\eta$ a small positive value controlling the precision of the approximation. The goal of the randomized power method is to compute a matrix $\bm{Q}\in\mathbb{R}^{n\times p}$, where $p=k+q$, such that the range of $\bm{Q}$ approximates the range of the top $k$ eigenvectors of $\bm{A}$, which are denoted $\bm{U}_{k}$. Specifically, we aim to ensure that:
$\|\bm{U}_{k}-\bm{Q}\bm{Q}^{\top}\bm{U}_{k}\|\leq\eta.$
The privacy-preserving randomized power method extends this approach to protect sensitive information in $\bm{A}$ during computation. Prior work (Hardt & Price, 2014; Balcan et al., 2016) show that under approximate Differential Privacy (DP) parameters $(\epsilon,\delta)$, $\bm{Q}$ can be computed using Algorithm 1, where the sensitivity $\Delta_{l}$ at iteration $l$ is bounded by $\Delta_{l}^{\text{prior}}\triangleq\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$. Here, $\Delta_{l}^{\text{prior}}$ represents an $\ell_{2}$-sensitivity bound that quantifies the impact of a single-element change in $\bm{A}$ on $\bm{A}\bm{X}^{l-1}$. Specifically, $\Delta_{l}\leq\Delta_{l}^{\text{prior}}$.
With this bound [^2] As defined in Fig. 3 of (Hardt & Price, 2014), Alg. 2 of (Balcan et al., 2016) and Theorem 6 of (Guo et al., 2021)., the standard deviation of the noise $\sigma_{l}$ associated to DP at iteration $l$ is given by:
$\sigma_{l}$ $=\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}$ (6)
$=\Delta_{l}^{\text{prior}}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}.$ (7)
As a first contribution, we refine this sensitivity analysis by introducing a significantly tighter bound for $\Delta_{l}$, reducing the noise required for DP and therefore allowing tighter convergence bounds (smaller $\eta$).
@@
1:  Input: Matrix $\bm{A}\in\mathbb{R}^{n\times n}$, number of iterations $L$, target rank $k$, iteration rank $p\geq k$, privacy parameters $\epsilon,\delta$.

### v3

@@ Abstract
The randomized power method has gained significant interest due to its simplicity and efficient handling of large-scale spectral analysis and recommendation tasks. However, its application to large datasets containing personal information (e.g., web interactions, search history, personal tastes) raises critical privacy problems. This paper addresses these issues by proposing enhanced privacy-preserving variants of the method. First, we propose a variant that reduces the amount of the noise required in current techniques to achieve Differential Privacy (DP). More precisely, we refine the privacy analysis so that the Gaussian noise variance no longer grows linearly with the target rank, achieving the same $(\varepsilon,\delta)$-DP guarantees with strictly less noise. Second, we adapt our method to a decentralized framework in which data is distributed among multiple users. The decentralized protocol strengthens privacy guarantees with no accuracy penalty and a low computational and communication overhead. Our results include the provision of tighter convergence bounds for both the centralized and decentralized versions, and an empirical comparison with previous work using real recommendation datasets.
@@ 1 Introduction
The randomized power method has emerged as an efficient and scalable tool for addressing large-scale linear algebra problems central to modern machine learning pipelines. By constructing an orthonormal basis for a matrix’s range in near-linear time, the method scales seamlessly to practical large datasets. Its reliance on simple matrix products ensures compatibility with sparse data representations and enables efficient parallelization and hardware acceleration on GPUs and distributed architectures.
Beyond its simplicity, the method provides strong approximation guarantees and accelerates a wide spectrum of applications. It has been used for principal component analysis (PCA) (Journée et al., 2010), singular value decomposition (SVD) (Halko et al., 2011), truncated eigen-decompositions (Yuan and Zhang, 2013), and matrix completion (Feng et al., 2018). Extensions have powered recommender systems (e.g., Twitter (Gupta et al., 2013), GF-CF (Shen et al., 2021), BSPM (Shen et al., 2021)), PageRank-style ranking (Ipsen and Wills, 2005), PDE solvers (Greengard and Rokhlin, 1997), or large-scale least-squares and linear-system solvers (Rokhlin and Tygert, 2008).
@@
However, as these methods are integrated into large-scale machine learning systems involving personal data, protecting user privacy is paramount. Unfortunately, the standard randomized power method does not inherently provide privacy guarantees. While its output might seem less sensitive than the input data, there is no formal guarantee against inference of private information embedded in the data.
@@
To address and quantify privacy leakages, Differential Privacy (DP) has emerged as a powerful framework that provides strong guarantees and mitigates potential privacy leaks of an algorithm, ensuring that the output of an algorithm reveals little about any individual record in the input. Several works have attempted to apply DP to the randomized power method. For example, Hardt and Price (2014); Balcan et al. (2016) developed centralized DP variants of the power method, whereas Wang and Xu (2020); Guo et al. (2024) investigated federated DP protocols that can be used when data is kept locally across multiple devices. Adjacent works explore DP versions of PCA in both centralized (Liu et al., 2022) and federated (Wang and Xu, 2020; Briguglio et al., 2023) settings and give optimal convergence bounds under distributional assumptions.
@@
Despite these advancements, existing approaches suffer from several limitations. First, their performance heavily depends on the number of singular vectors being computed (Hardt and Price, 2014; Balcan et al., 2016; Guo et al., 2024; Liu et al., 2022; Wang and Xu, 2020), which impacts both utility and privacy guarantees. Second, they are designed for centralized settings (Hardt and Price, 2014; Liu et al., 2022), where a trusted curator is assumed to hold the data. Moreover, some methods (Grammenos et al., 2020; Liu et al., 2022) make strong assumptions about the data distribution (e.g., sub-Gaussianity) which makes it harder to use these methods in practice. Some federated versions (Briguglio et al., 2023; Hartebrodt et al., 2024) claim to guarantee privacy due to the federated setting, but it has been shown that decentralization does not offer privacy by design (Geiping et al., 2020).
MOD-SuLQ (Chaudhuri et al., 2013) and its federated and streaming PCA variants (Grammenos et al., 2020) offer strong guarantees but are specifically tailored for reconstructing the top principal component ($k=1$). These methods add noise directly to the covariance matrix, rely on direct, computationally costly exact singular value decompositions (SVD) and are tailored for settings for which the number of samples largely exceeds the dimensionality ($n\gg d$). In contrast, the randomized power method iteratively adds noise directly to the approximation of the singular vectors themselves rather than the input matrix. This strategy reduces the dependence on the dimensionality of the original data, making it computationally more efficient and scalable. Memory-limited, streaming PCA methods such as those proposed by Mitliagkas et al. (2013) are optimized for sequential processing under memory constraints but lack privacy guarantees, requiring additional modifications.
Finally, no fully decentralized versions exist to our knowledge, making them unsuitable for decentralized environments (e.g., recommender systems and social networks), where data is partitioned across users/devices and communications are restricted to a predefined communication graph. The previously introduced DP federated versions (Balcan et al., 2016; Wang and Xu, 2020; Guo et al., 2024) use public channel communication and therefore require local DP, which hinders convergence.
@@ 1.1 Contributions
In this work, we propose a low-noise, privacy-preserving variant of the randomized power method whose approximation error remains bounded regardless of subspace dimension, significantly improving its applicability to practical datasets. We additionally propose a decentralized variant that strengthens privacy guarantees while maintaining the computational efficiency, and approximation guarantees of the centralized method. We summarize our contributions below:
  • An Improved Differentially Private Randomized Power Method. We introduce a Differentially Private variant of the randomized power method under a generalized adjacency model, making it more suitable for practical applications. Unlike prior works, which relied on loose sensitivity bounds extrapolated from one-dimensional vectors, we derive a significantly tighter sensitivity analysis directly applicable to the multidimensional case. We show theoretically with refined convergence bounds and empirically with a recommender system use-case [^1] We provide an anonymous code repository here. that this refinement considerably reduces the amount of noise and the approximation error for a given privacy budget.
  • Decentralized private power method. We propose a decentralized version of our algorithm, allowing settings for which data is distributed across multiple users or devices (e.g., recommender systems). This version is compatible with Secure Aggregation (Bell et al., 2020; Bonawitz et al., 2017; Kadhe et al., 2020), or fully decentralized private averaging approaches using correlated noise (Sabater et al., 2022; Allouah et al., 2024). Our decentralized approach preserves the effectiveness of the centralized version while incorporating Distributed Differential Privacy.
@@
  • Rigorous privacy guarantees. We present a new, more straightforward proof of a privacy result for the differentially private randomized power method. This proof addresses some mistakes in a privacy proof from Hardt and Roth (2012, 2013) that have been reproduced in several follow-up works (Hardt and Price, 2014; Balcan et al., 2016) and allows for a wider range of DP parameters.
The remainder of the paper is organized as follows: In Section 2 we review the necessary background and notation. Section 3 introduces our generalized adjacency notion, derives the improved sensitivity bound (Theorem 3.1), and presents the overall privacy proof (Theorem 3.2) together with runtime-dependent convergence guarantees (Theorem 3.3). In Section 4 we turn this into a fully runtime-independent bound (Theorem 4.1). Section 5 develops the decentralized variant (Algorithm 2), proves its equivalence to the centralized version (Theorem 5.2), and analyzes its communication and computation overhead. We empirically compare the methods on standard recommendation datasets in Section 6.1, and conclude in Section 7 with a discussion of limitations and future directions.
2 Background Material and Related Work
Matrix Norms and Notations.
@@ 2 Background Material and Related Work
For any matrix $\bm{X}$, the element-wise maximum norm is defined as $\|\bm{X}\|_{\max}=\max_{i,j}|\bm{X}_{ij}|$, where $\bm{X}_{ij}$ is the $(i,j)$-th element of $\bm{X}$. The $\ell_{2}$-norm is denoted as $\|\bm{X}\|_{2}$ and the Frobenius norm as $\|\bm{X}\|_{F}$. We use $\bm{X}_{j:}$ to denote the the $j$-th row of $\bm{X}$.
Eigenvalue Decomposition.
Let $\bm{A}\in\mathbb{R}^{n\times n}$ be a real-valued positive semi-definite matrix, where $n$ is a positive integer. The eigenvalue decomposition of $\bm{A}$ is given by $\bm{A}=\bm{U}\bm{\Lambda}\bm{U}^{\top},$ where $\bm{U}\in\mathbb{R}^{n\times n}$ is a matrix of eigenvectors and $\bm{\Lambda}\in\mathbb{R}^{n\times n}$ is a diagonal matrix containing corresponding eigenvalues.
QR Decomposition.
We will use the matrix QR decomposition, obtained using the Gram-Schmidt procedure. Given a matrix $\bm{X}\in\mathbb{R}^{n\times p}$, the QR decomposition factorizes it as $\bm{X}=\bm{Q}\bm{R}$, where $\bm{Q}\in\mathbb{R}^{n\times p}$ is an orthonormal matrix (i.e., $\bm{Q}^{\top}\bm{Q}=\bm{I}$) and $\bm{R}\in\mathbb{R}^{p\times p}$ is an upper triangular matrix.
Gaussian Random Matrices.
We denote by $\mathcal{N}(\mu,\sigma^{2})^{n\times p}$ a $(n\times p)$ random matrix where each element is an independent and identically distributed (i.i.d.) random variable according to a Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$.
Coherence Measures of a Matrix.
We define below two coherence measures for the matrix $\bm{A}$, which will allow us to give convergence bounds for our method.
@@
  • The $\mu_{1}$-coherence of $\bm{A}$ is the maximum row $\ell_{2}$-norm of its eigenvectors, defined as $\mu_{1}(\bm{A})=\|\bm{U}\|_{2,\infty}=\max_{i}(\|\bm{U}_{i:}\|_{2})$.
2.1 Differential privacy
With positive integers $n$ and $m$ specifying the matrix dimensions, let $D_{1}\in\mathbb{R}^{n\times m}$ and $D_{2}\in\mathbb{R}^{n\times m}$ be two matrices representing two datasets embedding sensitive information. $D_{1}$ and $D_{2}$ are said to be adjacent ($D_{1}\sim D_{2}$) if they differ on one sensitive element of the dataset. For example, in a recommender system application, $D_{1}$ and $D_{2}$ can be binary user-item interaction matrices and a sensitive element of the dataset can be an entry in the matrix, corresponding to a user-item interaction.
A randomized algorithm $\mathcal{M}$ provides ($\epsilon,\delta$)-Differential Privacy (DP) if for all adjacent datasets $D_{1}$ and $D_{2}$, and for all measurable subsets $S\subseteq\text{Range}(\mathcal{M})$, the following holds:
@@ 2.1 Differential privacy
where $\epsilon$ is a small positive scalar representing the privacy loss (smaller values indicate stronger privacy guarantees), and $\delta$ is a (typically negligible) probability that the privacy guarantee fails.
Let $f:\mathbb{R}^{n\times m}\to\mathbb{R}^{d}$ be a query function associated with a mechanism $\mathcal{M}$. DP guarantees of mechanisms are defined using the sensitivity of $f$. In our contribution we use the $\ell_{2}$-sensitivity, denoted by $\Delta_{2}(f)$ or $\Delta_{2}$ and defined as $\Delta_{2}=\max_{D_{1}\sim D_{2}}\|f(D_{1})-f(D_{2})\|_{2}\,.$
2.2 Privacy-Preserving Randomized Power Method
@@
Let $k$ be the target rank, let $b>0$ be a small positive integer, and let $\eta>0$ be an approximation tolerance. Given a positive semi-definite (PSD) matrix $\bm{A}\in\mathbb{R}^{n\times n}$, the aim of the randomized power method is to construct a matrix $\bm{Q}\in\mathbb{R}^{n\times p}$, where $p=k+b$, whose column space approximates that of the top $k$ eigenvectors of $\bm{A}$, i.e., $\bm{U}_{k}\in\mathbb{R}^{n\times k}$. Specifically, it aims to satisfy
$\bigl{\|}\bm{U}_{k}-\bm{Q}\bm{Q}^{\top}\bm{U}_{k}\bigr{\|}\leq\eta.$ (2)
To protect sensitive information in $\bm{A}$ while computing $\bm{Q}$, prior work (Hardt and Price, 2014; Balcan et al., 2016) shows that the randomized power method can be implemented as Algorithm 1 for it to satisfy $(\epsilon,\delta)$-Differential Privacy with adjacency defined as a single element change in $\bm{A}$. The corresponding adjacency notion is defined in more detail in Equation (6).
To compute the standard deviation of the noise required to satisfy $(\epsilon,\delta)$-DP, one needs to bound the $\ell_{2}$-sensitivity $\Delta_{l}$ of the computations at each iteration of the algorithm. Prior works (Hardt and Price, 2014; Balcan et al., 2016) use the estimate
$\Delta_{l}^{\text{prior}}\triangleq\sqrt{p}\|\bm{X}^{l}\|_{\max},$ (3)
which upper-bounds the true sensitivity, i.e., $\Delta_{l}\leq\Delta_{l}^{\text{prior}}$. Here, $\Delta_{l}$ bounds the change on $\bm{A}\bm{X}^{l-1}$ under a single-element perturbation in $\bm{A}$.
Using this bound [^2] See Fig. 3 of Hardt and Price (2014), Alg. 2 of Balcan et al. (2016), and Theorem 6 of Guo et al. (2024)., the standard deviation $\sigma_{l}$ of the Gaussian noise added at iteration $l$ to achieve $(\epsilon,\delta)$-DP is $\sigma_{l}=\sqrt{p}\,\bigl{\|}\bm{X}^{l}\bigr{\|}_{\max}\epsilon^{-1}\sqrt{4L\ln(1/\delta)}\,,$ where $L$ is the total number of power iterations.
@@
1:  Input: Matrix $\bm{A}\in\mathbb{R}^{n\times n}$, number of iterations $L$, target rank $k$, iteration rank $p\geq k$, privacy parameters $\epsilon,\delta$